If the mom is nine years times her daughter's age after 4 years she will be 5 years times her daughter's age, what is the daughter's and mom's age?
Answer by ewatrrr(24785)


Hi,
mom is nine years times her daughter's age
Let x and 9x represent daughter's and mother's ages NOW respectivley
Write as we Read***
(9x+4) = 5(x + 4)
Solving for x
9x + 4 = 5x + 20
4x = 16
x = 4, is the daughter's age. Mother is 36 years old
CHECKING our Answer***
40 = 5*8